### v1

# Evaluating Time-Dependent Methods and Seasonal Effects in Code Technical Debt Prediction

Abstract
Background. Code Technical Debt prediction has become a popular research niche in recent software engineering literature. Technical Debt is an important metric in software projects as it measures the effort professionals need to clean the code. Therefore, predicting its future behaviour becomes a crucial task. However, there is no well-defined and consistent approach which is able to completely capture the features that impact on the evolution of Code Technical Debt.
Objective. The goal of this study is to evaluate the impact of considering time-dependent techniques as well as seasonal effects in temporal data in the prediction performance within the context of Code Technical Debt. The study adopts already existing, yet not extensively adopted, time-dependent prediction techniques and compares their prediction performance to that of commonly used Machine Learning models. Further, the study strengthens the evaluation of time-dependent methods by extending the analysis to capture the impact of seasonality in Code Technical Debt data.
Methods. We trained 11 prediction models for using the commit history of 31 open-source projects developed with Java. We predicted the future observations of the SQALE index to evaluate their predictive performance.
Results. Our study confirms the positive impact of considering time-dependent techniques. The adopted multivariate time series analysis model ARIMAX overcame the rest of the adopted models. Incorporating seasonal effects led to an enhancement in the predictive performance of the adopted time-dependent techniques. However, the impact of this effect was found to be relatively modest.
Conclusions. The findings of this study corroborate our position in favour of implementing techniques that capture the existing time dependence within historical data of software metrics, specifically in the context of this study, namely, Code Technical Debt. This necessitates the utilisation of techniques that can effectively address this evidence.
keywords:
Technical Debt, Software Quality Mining Software Repositories, Empirical Software Engineering, Time Series Analysis1 Introduction
In Software Engineering (SE), and especially in Mining Software Repository (MSR) studies, researchers often investigate the relationships among different variables collected from the history of software projects. As an example, researchers have investigated correlations between two variables such as code smells [1, 2, 3], their trend over time [4], and the impact of different software qualities [5, 6, 7]. However, a large number of these studies have omitted the potential impact hidden on the temporal dependency of the analyzed variables and the possible threats related to the usage of statistical techniques not designed for analyzing temporally dependent data [8]. As an example, the introduction of a technical issue in a commit heavily depends on the code that was present in the repository before the commit. Still, most studies have not considered this aspect, mainly due to a lack of clear guidelines on the statistical techniques that can be used in this context. Saarimäki et al. [8] highlighted three main issues in previous studies: 1) discarding the temporal nature of the commits, 2) assuming independence of data, and 3) mixing projects of different sizes, where big projects overwhelm small ones. To overcome these issues, they proposed to analyze dependent data in the SE field, considering the dependency through the time effect [8]. Within the provided study context, we leverage the analysis of the Code Technical Debt (TD) prediction to evaluate the impact of the time-dependence factor. Code TD is an important metric in software projects as it measures the effort professionals need to put into making the code clean. Therefore, Code TD denotes time dependence on past mistakes. Consequently, data analysis techniques that assume this dependence should be considered.
In this direction, the research community traditionally used Machine Learning (ML) models [9, 10]. Similarly, we identified two relevant works that applied Time-Series Analysis (TSA) to predict TD [11, 12] comparing univariate and multivariate TSA models. The obtained results were promising in terms of prediction performance. Hence, we aim to analyze the robustness of TSA techniques predicting time-dependent variables such as Code TD by comparing them with traditional ML methods. Therefore, we aim to identify existing potential variables that can help explain the behavior of Code TD.
We designed and conducted an empirical study to evaluate the Code TD prediction performance of different TSA models and a set of time-agnostic ML models commonly used in the literature, which do not take into consideration the temporal factor of the data. Furthermore, we assessed the impact of the seasonality factor in the observed data. For that, we adopted the seasonally adjusted version of the TSA models already used in SE literature and measured their performance to quantify the impact of addressing the seasonality factor. As the adopted potential variable on the description of the Code TD, we considered the SQALE index (also called TD) computed by SonarQube (SQ) to depict TD. Specifically, we adopted two different multivariate applications of the Autoregressive Integrated Moving Average (ARIMA) model [13] as well as their modified counterparts which control the seasonality component of the data, and seven ML models to perform the confronted evaluation (four linear and three non-linear). Additionally, we implemented a backward variable selection approach [14, 15] to provide the models with the best variable combination to help improve their efficiency.
Our results revealed a robust superiority obtained by the multivariate TSA models and demonstrated a slight improvement in the predictive performance when adjusting the models for capturing the seasonality effect. Especially, the Seasonally adjusted Autoregressive Integrated Moving Average with extra regressors (SARIMAX) and its unadjusted counterpart ARIMAX resulted in being the best prediction models among the methods considered in this study, both the mentioned methods being the multivariate application of the commonly known ARIMA model. However, the ARIMAX model demonstrated to be the model robustly overtaking the rest of ML algorithms in terms of predictive performance.
Therefore, the main contributions of this paper are:
An improvement to the current knowledge of Code TD prediction with an emphasis on investigating the importance of including time-dependence factors in the prediction.
An additional contribution to the scientific community emphasizing the potential benefits of considering a monitored seasonal Code TD measurement for subsequent improvement on the control of Code TD in software projects.
A large-scale comparison among the time-dependent prediction models and ML prediction models already adopted in the literature.
Paper structure
Section 2 describes the background our paper is based on. In Section 3, we present the empirical study design. In Section 4, we present the obtained results, and Section 5 shows the derived discussion based on them. Section 6 describes the identified threats to the validity of our study. Section 7 presents related work, and Section 8 divulges the conclusions of this study and outlines the potential future work.
2 Background
SQ is one of the most common static analysis tools adopted both in academia [16, 17] and in industry [18].
It aims to help developers create more clean, secure, maintainable, readable, and modular code. SQ evaluates the code by testing it against a predefined set of rules. The rules are considered coding standards and breaking them is deemed undesirable. Therefore, each time a piece of code violates a rule, SQ generates an issue presented to the developers.
SQ can be accessed as a service through the sonarcloud.io platform or downloaded and run on a private server. The tool supports 29 different programming languages including Java, Python, C++, and JavaScript.111https://docs.SQ.org/9.6/analyzing-source-code/languages/overview/ It performs various calculations, including measuring metrics like lines of code and code complexity. It also establishes specific thresholds known as “quality gates” for each metric and rule.
In this study, the selected dataset (Section 3.2) analyzed the projects with SQ version 7.5.
This version includes three rule categories:
Reliability rules named as bugs that create an issue that “represents something wrong in the code” and that “will soon be reflected in a bug in the code.”
Maintainability rules named as code smells that decrease code readability and code modifiability. It is important to note that the term “code smells” adopted in SQ contains only a subset of the well-known code smells defined by Fowler et al. [19].
Security rules named as Vulnerability that creates a problem that impacts the application’s security.
SQ internally uses the SQALE methodology [20] to compute the technical debt ratio (TD ratio) and to classify the project to an SQALE rating considering the time to fix all open issues classified as code smell. Moreover, SQ computes other two types of debt called reliability remediation effort and security remediation effort, based on respectively on the time to fix all the open issues classified as bugs and as vulnerabilities.
3 The Empirical Study Design
We now describe our study reporting the goal and research questions, the context, data collection, and data analysis. We designed our experimental study based on the guidelines defined by Wohlin et al. [21].
3.1 Goal and Research Questions
We formalized the goal of this study according to the GQM approach [22] as follows:
Analyze time-dependence and seasonality effect factors,
for the purpose of comparing Code TD prediction,
with respect to performance,
from the point of view of developers,
in the context of open source software.
Based on this goal, our formulated Research Questions (RQs) are presented as follows:
RQ1. Which multivariate time-dependent method predicts Code TD with the highest performance?
The defined RQ1 compares two different multivariate TSA approaches already implemented in the SE literature to determine which approach can predict better Code TD. Evaluating the prediction performance of the latest models used to date that control the time-dependence factor within the data enables us to determine which of these models should be considered by practitioners when deliberating which prediction techniques to use when predicting Code TD. To address RQ1, we collected as Code TD the SQALE index metric computed by SQ (as explained in Section 2), as well as Code Smell rule violations generated by SQ as potential explainers of the SQALE index’s behavior.
The first approach, referred to as ARIMA + LM, combines the use of multiple ARIMA models and linear regression models (LM) [12]. This approach utilizes a univariate ARIMA model for the prediction of each independent variable and utilizes the predicted values inside a regression model to predict future values of the dependent variable. The second approach, referred to as ARIMAX consists of constructing a multivariate ARIMA model for predicting values of the dependent variable [11]. In the ARIMAX model, the history of the dependent variable as well as the past values of the considered independent variables are used as features to train the model and provide future predictions of the dependent variable.
Our prior assumption is that the ARIMAX model performs better than the ARIMA + LM models in explaining the future evolution of the dependent variable. We assume that combining the independent variables with the past values of the dependent variable has more information than separately predicting the values of the independent variables to predict the dependent variable through a regression model.
The goal of this study is not only to establish the applicability of the most efficient time-dependent techniques for Code TD prediction but also to evaluate how the adopted techniques perform against other prediction techniques commonly used in the SE literature. These techniques are mostly ML-based algorithms that do not consider the temporal factor when providing future predictions. Therefore, we formulate our RQ2:
RQ2. How accurate is the prediction performance of a multivariate time-dependent approach for Code TD prediction compared to that of an ML algorithm?
To assess the applicability of a time-dependent approach for Code TD prediction (RQ2), in our case multivariate TSA, we need to compare its performance with other approaches that do not take into consideration the time dependency of the data. In this study, the comparison is made with several ML models.
Similarly, as for RQ1, our prior assumption is that the prediction performance of multivariate TSA methods for Code TD prediction overcomes that of ML models. Moreover, TSA methods are based on temporally ordered serialized data, which helps the models capture potential temporal factors, for instance, the seasonality of the observation [23]. Models can be adjusted to these factors to observe their impact on the predicted results. Therefore, only comparing already used TSA methods is not sufficient and we formulate a third research question as follows:
RQ3. Does addressing seasonality improve the prediction performance of multivariate time-dependent approaches for Code TD prediction?
As explained, RQ3 expands RQ1. For that, given the modification in the models to handle seasonality patterns, RQ3 compares the new models with the former ones in terms of Code TD prediction and therefore evaluates the impact of seasonality in the prediction performance.
Our prior assumption is that SARIMAX, SARIMA + LM approaches and the seasonally adjusted models of the previously defined TSA model counterparts perform better than their former model format which did not handle seasonality patterns. Further, we leverage the comparison proposed in RQ1, thus we assume that using the multivariate SARIMAX model is more informative than the SARIMA+LM model, given the extension of both TSA models to capture the impact of the seasonality effects.
Similarly, evaluating the seasonality adjustment in the prediction performance compared to the selected ML algorithms becomes paramount to providing a complete contribution to the SE community through this study. Therefore, we ask:
RQ4. How accurate is the prediction performance of a seasonally adjusted multivariate time-dependent approach for Code TD prediction compared to that of an ML algorithm?
Our prior assumption is that the additional modification implemented in the multivariate TSA models regarding seasonality control enables their Code TD prediction performance to be better than that of ML models.
For all RQs, the models were evaluated using the following set of performance metrics: Mean Absolute Percentage Error (MAPE), Mean Absolute Error (MAE), and Root Mean Square Error (RMSE) which will be further described in Section 3.5.5.
3.2 Context
As the context of our study, we consider the Technical Debt Dataset [24] version 2.0 adopted by Saarimäki et al. [8]. The data set contains 31 Java projects from the Apache Software Foundation (ASF) repository.222http://apache.org The projects in the data set were selected based on “criterion sampling” [25], that fulfil all of the following criteria: developed in Java, older than three years, more than 500 commits and 100 classes, and usage of an issue tracking system with at least 100 issues reported. The projects were selected, maximizing their diversity and representation by considering a comparable number of projects concerning project age, size, and domain. The projects can be considered mature, due to the strict review and inclusion process required by the ASF. Moreover, the included projects regularly review their code and follow a strict quality process.333https://incubator.apache.org/policy/process.html Full details of the data can be found in the online repository of the data set.444https://github.com/clowee/The-Technical-Debt-Dataset We adopted this dataset since all its projects were already analyzed with SQ and it contains the data related to the SQALE Index.
3.3 Variables
In this section, we describe the dependent and independent variables used in the adopted models for Code TD prediction:
Dependent variable: We considered the metric SQALE index computed by SQ. The SQALE index denotes the efforts to fix all Code Smell rules violated in the code and is measured in minutes. SQ denotes TD using the total value of the SQALE index. The metric is continuous and, therefore, is suitable for the TSA models adopted in this study. Further, it provides a clear definition of Code TD to answer the goal of our study.
Independent variables: We considered the Code Smell rule violations detected by SQ. Code smell rule violations describe the number of violations for each SQ rule of code smell type. The TD dataset has issues from 205 different rules of type code smell and each rule is included as a separate independent variable in the study. The nature of the collected issues is discrete, as it is founded on the count of the issues detected in each SQ analysis for each of the Code Smell types.
3.4 Data Collection
This study included 31 projects from the previously described Technical Debt Dataset. For this study, we used all commits analyzed by SQ and extracted their commit information, SQALE index, and code smell rule violations reported by SQ.
3.4.1 Data preprocessing: Feature selection
The collected data contains 205 potential independent variables. Such a large number could have a negative impact on the models due to the so-called “curse of dimensionality”: including a large number of independent variables may drastically reduce the predictive power of a model. To identify a set of the most explicative independent variables for our dependent variable, we undertook a feature selection process that utilized well-known feature selection techniques. The aim was to categorize the potential independent variables according to their importance. These techniques are briefly described next.
Variance Thresholding: A method utilizing the variance or spread of a variable in a data set for selecting model-independent variables. In a nutshell, this method calculates the variance of all the potential independent variables across the data sets and removes those with low variance as they are less informative and do not have much predictive power [26].
Zero percentage method: This method eliminates predominantly empty independent variables, i.e. the percentage of emptiness is higher than a predefined threshold, and is therefore poorly informative. It is used with count data [27].
Feature importance: This method determines the importance of each variable in a data set using a Random Forest (RF) algorithm. RF is an ensemble learning technique consisting of multiple decision trees, in this case, used for prediction. The measure of importance is based on the reduction in node impurities due to the splitting of the variable. Averaged over all trees, the node impurity measure of all the independent variables is calculated by the residual sum of squares. Thus, this method excludes variables with low feature importance [28].
Correlation analysis: The method evaluates the degree of linear association between the independent variables and the dependent variable. Given the high skewness illustrated in Figure 1, and therefore lack of normality in our variables [29], we measure the correlational association through the non-parametric Spearman’s rank correlation coefficient [30] for each independent variable. Thus, variables with a poor informative contribution for predictive modelling are discarded [31].
Understandably, one could expect that each of the selected feature selection techniques would provide different results on the initially collected issue type variables. Therefore, we ranked the independent variables in terms of their resulting importance score for each of the feature importance techniques, thus resulting in four different rankings. We used the Interquartile Range (IQR) method to select the uppermost quartile in the distribution of the rankings based on the importance scores. In a nutshell, given a set of observations, the IQR method builds underlying the distribution from the observations and selects the subset of the different quartiles in the distribution based on the goal of the analysis [32]. In our scenario, we were interested in selecting the most informative independent variables of our set of variables; therefore, we selected the quartile with the variables denoting the highest feature importance score. Additionally, for the sake of consistency in the final set of independent variables, we selected the subset of variables that showed importance in each of the four fetched quartiles. Consequently, the feature selection process resulted in a subset of 15 independent variables that were specifically selected to predict the SQALE index (dependent variable). Table 1 and Figure 1 provide a graphical and quantitative portrayal of the independent variables.
3.4.2 Data preprocessing: Transforming the raw data into time series data
The time-dependent approaches used in this paper require using periodically measured time-serialized data. Given that commits are not periodical, we relied on two criteria to generate the time series. First, we focused on constructing a sufficiently long time series for each of the projects (study subjects) to provide the models with sufficient data to perform the prediction. Second, we investigated different periodicity levels for the observations in the time series. In this way, we could observe the temporal nature of the data and find the most suitable time frame between observations.
Based on the described criteria we generated time series data at biweekly and monthly periodicity levels. The number of data points for each project on each periodicity level is presented in Table 2. TSA models require at least 24 observations to detect existing time-dependence factors such as the temporal trend, and consequently to be trained [23]. Therefore, increasing the time window size would reduce the number of observations for the collected projects, and require excluding projects from the study.
Version control data is stochastically ordered by nature555https://www.git-scm.com/docs/git-commit, which means that some time periods (of weeks or months in our study) might not have commits despite otherwise being an active project. Consequently, serializing such data can result in missing data for some time points. Linear interpolation was used to fill in time points that lacked observations in the generated time series. This method presupposes a linear correlation between neighboring data points and calculates the values of the missing data points using the known values surrounding them [33].
Consequently, the analysis was conducted using three different samples of the data: monthly time series data, biweekly time series data, and original commit data, which is chronologically ordered but not serialized in periods, and is only used with the ML models. As a potential threat to the validity of this study, we address the interpolation of artificially generated observations in Section 6. Similarly, we provide suggestions for practitioners on the remediation approach to avoid missing data in Section 4.1 to help practitioners with generating higher-quality data.
| Variable | Mean | Standard deviation | Min | Lower quartile | Median value | Upper quartile | Max | Skewness |
| SQALE index | 137,726 | 169,927 | 0 | 24,378 | 72,981 | 182,132 | 701,914 | 1.87 |
| S1213 | 740 | 1,350 | 0 | 38 | 152 | 579 | 6,106 | 2.22 |
| RedundantThrows DeclarationCheck | 345 | 509 | 0 | 15 | 113 | 340 | 2,200 | 1.91 |
| S00117 | 1,082 | 2,380 | 0 | 7 | 55 | 383 | 10,395 | 2.3 |
| S00122 | 402 | 658 | 0 | 1 | 67 | 585 | 4,119 | 2.55 |
| S1488 | 88 | 163 | 0 | 4 | 20 | 81 | 1,196 | 2.85 |
| S1905 | 74 | 148 | 0 | 1 | 12 | 44 | 575 | 2.26 |
| UselessImportCheck | 630 | 1,402 | 0 | 7 | 42 | 151 | 7,928 | 2.58 |
| DuplicatedBlocks | 333 | 510 | 0 | 40 | 142 | 305 | 3,099 | 2.52 |
| S1226 | 212 | 293 | 0 | 18 | 62 | 330 | 1,479 | 1.91 |
| S00112 | 473 | 848 | 0 | 58 | 183 | 374 | 4,987 | 3.01 |
| S1155 | 86 | 169 | 0 | 2 | 12 | 67 | 933 | 2.49 |
| S00108 | 97 | 128 | 0 | 4 | 62 | 127 | 739 | 1.95 |
| S1151 | 472 | 1,007 | 0 | 2 | 26 | 200 | 3,699 | 2.25 |
| S1132 | 289 | 481 | 0 | 20 | 57 | 436 | 3,587 | 2.94 |
| S1481 | 132 | 274 | 0 | 5 | 18 | 71 | 1,600 | 2.75 |
3.5 Data Analysis
In this section, we focus on the designed data analysis based on the definition of the selected models, and explain the comparison of the seasonally adjusted models against their root model counterparts, their model parameter tuning structure, and the ML models adopted. The selected time-dependent models are based on multivariate TSA modelling, a field of statistical modelling commonly used in diverse fields such as econometrics and weather forecasting for instance [34]. The multivariate nature of the models allows us to consider information about additional independent variables that can assist the model in the prediction process.
3.5.1 Setting the stage for seasonally unadjusted TSA model prediction
In order to address RQ1, we examine two distinct methods to predict Code TD utilizing TSA models, ARIMAX and ARIMA+LM. As version control data is inherently non-stationary, i.e. denote a trend across time, the two methods originate from the Box-Jenkins model. better known as Auto-Regressive Integrated Moving Average (ARIMA) model. It is worth noticing that a time series of data is stationary when its predictions do not depend on factors such as trend or seasonality existing in the real data [35]. The ARIMA model is a modified version of the Autoregressive Moving Average (ARMA) model used for modelling non-stationary time series [36]. Like the ARMA model, ARIMA determines the best autoregressive parameter for the variable to be predicted (AR step) and also identifies the optimal moving average parameter (MA) [34]. In addition, ARIMA executes the integration step (I), which determines the level of differencing required for the time series of the variable in question to achieve stationarity. The methods considered are:
ARIMAX: The approach used in [11] was an ARIMA model with a set of independent variables included in the model, commonly referred to as ARIMAX. This model offers a further multivariate approach to the ARIMA model by introducing additional independent variables into the model and, therefore, helping to explain the evolution of the SQALE index. This enables future SQALE index values to be explained by both its past values and those of the independent variables.
ARIMA + LM: An ensemble technique previously adopted in [12] that comprises constructing a univariate ARIMA model for each independent variable in the model. Each ARIMA model considers the past values of their respective model variable as model parameters and, therefore, independently predicts the future values of each model variable. The additional step in this methodology relies on performing Linear Regression (LM) considering each of the new prediction lag values to predict the new values for the SQALE index variable.
3.5.2 Setting the stage for Seasonally Adjusted TSA model prediction
In contrast with the previously presented ARIMAX and ARIMA+LM models, the seasonally adjusted models capture the impact of the seasonality pattern in the data while maintaining the model nature of the previously defined TSA model counterparts. Therefore, the analyzed seasonally adjusted methods are:
SARIMAX: The Seasonally adjusted Auto-Regressive Integrated Moving Average model stands as the seasonally adjusted extension of the ARIMAX model. It addresses the seasonality of the data within the model training through the model parameter tuning [13]. Further information regarding the model parameter tuning in the TSA model is provided later in Section 3.5.3. Similarly, the SARIMAX model provides a multivariate approach compared to its univariate counterpart, SARIMA, by including independent variables in the model. This enables future SQALE index values to be explained by both its past values and those of the included independent variables.
SARIMA + LM: The approach stands as the combination of the univariate Seasonally adjusted Auto-Regressive Integrated Moving Average (SARIMA) model combined with a Linear Regression model (LM). Based on the seasonality adjustment performed in the approach used in [12], a SARIMA model is built per each of the considered independent variables in the study, and thus each model independently predicts the future values of a single variable accordingly. The additional step in this methodology relies on performing Linear Regression (LM) considering each new prediction lag value of the independent variables to predict the new values for the SQALE index variable.
3.5.3 Setting the model parameter order in TSA models
Similarly as in other families of predictive analysis models, for instance, Linear Models or Generalized Linear Models [29], TSA models are founded on a set of model parameters which are estimated after the model is trained. As mentioned earlier in this section, in this study we adopt as prediction models derivations from the Box-Jenkins ARIMA model [13]. Models founded on the ARIMA model follow the same model parameter structure, commonly depicted as follows:
where p stands as the number of steps that a model needs to go back into the past history data to predict the next value i.e, autoregression (AR) level, q denotes the number of moving averages required (MA) and captures the short-term random fluctuations in the data, and d defines the differencing level required to achieve stationarity in the trend observed within the data and therefore be able to make predictions [37]. Models must be adjusted to these factors in order to reach stationarity on the trained data. We provide a detailed description of how parameter tuning is undergone with ARIMA-based models in the Appendix.
Similarly, models capturing seasonal patterns such as SARIMA are presented with an extended set of parameters
When the model is adapted to capture the seasonal patterns of the data in models such as SARIMAX and SARIMA+LM, the parameters P, D, Q are calculated as well. These capture the impact of seasonality in the model parameters explained before, which the model tunes based on the specified m cycle level in the data. In our study, we consider monthly and biweekly cycles as previously defined. Thus, the adjusted models provide an additional layer in the model parameter tuning where the seasonality component is treated.
In this study, the presented model parameters are tuned when the TSA models are fitted with the training data. To determine the optimal combination of the model parameters, we conduct an iterative process in which we search for the best model parameter combination for each of the project’s time series data. Thus, for different combinations of and parameter values, we utilize the Augmented Dickey-Fuller test [38] to determine the optimal value d of the differencing model parameter. Similarly, when the seasonality effect is adjusted, the Canova-Hansen test [39] is used to determine the optimal order of seasonal differencing as well, along with different combinations of and . To execute this iterative process, we adopt the auto arima algorithm [40, 41] which defines the potential range of parameter values based on the fitted data.
Thus, based on the temporal patterns of the SQALE index, the computed different model combinations of the defined parameters are assessed through Akaike Information Criterion [42] (AIC) and Bayesian Information Criterion [43] (BIC) following the state-of-the-art in statistical modelling [29]. We provide the theoretical description of the presented model parameters in Appendix Appendix: Parameter tuning of the Box-Jenkings models. Moreover, we provide the implemented logic in the shared online appendix (see Section 3.7).
3.5.4 Setting the stage for Machine Learning models
In RQ2 and RQ4, we compare the prediction performance of TSA approaches with ML algorithms. Therefore, we also investigate the ability of ML models to predict Code TD by applying a collection of linear and non-linear ML algorithms with the collected data and compare the results with the time-dependent approaches mentioned before.
The ML methods used for the comparison are selected to provide a wide predictive perspective, as each of them emphasizes different aspects of the data. Likewise, the chosen ML algorithms have been extensively used in the recent literature for their ability to predict software quality features, like TD [12, 9, 10, 44, 45].
The included ML models assuming linearity (L) in the data are the following.
Multiple Linear Regression (MLR) is a commonly used and powerful statistical model for analyzing the relationship between a continuous dependent variable and a set of independent explanatory variables. Through this relationship, the model can provide insights on how changes in the independent variables affect the value of the response variable while maintaining the linearity assumption [46].
Stochastic Gradient Descent (SGD) algorithm relies on the iterative decrease of prediction loss by changing the model parameters. Precisely, for each iteration, the SGD algorithm randomly selects a training set to bring randomness into the optimization problem, and the algorithm continues optimizing the result until it converges [47].
Lasso regression (L1) is a linear regression technique specialized in high dimensional explanatory variable sets. This technique performs an optimization problem where model-independent variables get penalized if they are not important for the model. Thus, through model simplification, L1 achieves linear regression with the most relevant explanatory variables [48].
Ridge regression (L2) is another linear regression technique that, through a regularization term, avoids overfitting the model. The L2 technique brings a penalization for the given model parameters that denote extreme values and thus stabilizes the model which results in advantages for applications where the number of model-independent variables is high [49].
The selected ML models assuming non-linearity (NL) are:
Support Vector Machine (SVM) is a well-established ML algorithm used for classification and regression tasks. The regression model aims to find the most optimal hyperplane that considers the characteristics of the existing explanatory variables for explaining the values of the output response variable. For this, the model performs a minimization process of the deviation between the real value and the predicted value [50]. We considered using the so-called Gaussian kernel or Radial Basis Function as the function to build the model to get the perspective of a non-linear approach.
Extreme Gradient Boost (XGB) is a ML algorithm part of the gradient boosting method family. Based on the ensemble learning methodology, the XGB algorithm employs multiple decision trees in a greedy format, where each of the subsequent trees corrects or refines the wrong results from the previous tree. Through this technique, the algorithm defines weights for the included independent variables that require a higher emphasis [51].
Random Forest regression (RF) also belongs to the ensemble learning family, and it is built upon the basis of the ensemble of decision trees. Based on random sampling, the RF algorithm performs multiple decision trees and the output is the aggregated value from the results obtained in the performed trees [52].
3.5.5 Performance metrics
We used four performance metrics to compare and evaluate the prediction approaches and models.
The first performance metric is the Mean Absolute Percentage Error (MAPE). The MAPE is a statistical metric commonly used for measuring prediction accuracy. It quantifies the magnitude of errors between the predicted values and the actual observed values, calculating the mean of the absolute value of the prediction error. Its formula is:
where is the number of observations, is the i-th actual observed value and is the i-th predicted value. As shown by the formula, MAPE is represented as a percentage metric, where a smaller percentage of error means better predictive performance and the opposite with a higher percentage.
MAPE has limitations when the actual observations are small or close to zero, thus introducing bias into the model training. Therefore, we selected alternative accuracy measurement statistics to cover this issue. One of them, and the second metric to present is the Mean Absolute Error (MAE), which is characterized by measuring the average magnitude from the absolute value of prediction errors. Its equation is:
where the equation variables follow the same notations as the ones addressed for MAPE. MAE expresses the prediction performance error in absolute value; therefore, a smaller result signifies better predictive performance, while higher results represent poorer performance.
We adopt as the third performance metric Root Mean Squared Error (RMSE) which captures the error value in the same value unit as the variable being predicted. Its formula is defined by
where the equation variables follow the same notation as the ones explained previously. Although the definitions have value unit differences, a low error value indicates a high predictive performance in RMSE and similarly, a high error value depicts a poor predictive performance. We provide further results on AIC and BIC model selection criteria in the provided online appendix (see Section 3.7).
3.5.6 Performance evaluation
The next step was to train and evaluate the performance of the selected prediction approaches using the three data sets explained in Section 3.4.2. The models were trained and tested using the Walk-Forward Train-Test validation technique [53]. This technique was chosen as it respects the temporal nature of the data and does not rely on randomness, yielding a fair comparison between methods. The Walk-Forward Train-Test technique trains and tests a model greedily and iteratively. Each iteration trains a model, predicts the value of the next data point of the time series, and compares the prediction against the real value. The real value is added consequently to the train set and the model is trained again with the newly extended train set data. This cycle is continued until all test data points are predicted. In our study, the models initially start the training with 80% of the data as training data and subsequently predict and test every new data point. Figure 2 shows a graphical representation of the approach, where i denotes the new data point in each iteration. The described model training and testing phases performed in the performance evaluation stage of this study were implemented equally in each of the adopted models, with the same testing set of data, for the sake of fairness in the predictive performance comparison.
3.6 Model Execution
In this section, we emphasized the description of the model execution process of the defined TSA models to facilitate practitioners’ understanding of the results presented in Section 4 and encourage the use of the proposed models in the SE community. Nonetheless, the prediction process described in Section 3.5.6 is similarly followed for all projects with each of the considered prediction models using different data sets. Note that the ML models are tested with all three data sets, but the time-dependent models only use the biweekly and monthly time series data. After the prediction process is concluded, the results are visualized, and therefore organized, by aggregating the results provided by each model with all the collected projects for the defined three data sets (raw data, biweekly data, and monthly data). The aggregation of the results per dataset for each model combination is done by taking the average of the performances for the different projects in the respective data set as displayed in Tables 4 and 5 for instance.
3.6.1 Execution process for ARIMAX & SARIMAX models
Within the context of the modelling process, both ARIMAX and SARIMAX models are similarly built in their initial stage. The last one stands as an adjusted version of the former counterpart. Therefore, for each project, we built both an ARIMAX and SARIMAX model. Given the existing lack of normality in the distribution of the independent variables (see Table 1 and Figure 1) we log-transformed their observations in order to standardize the independent variable observations. Variable transformation is a common preprocessing technique used in the field of predictive analytics due to its benefits in reducing the model complexity and augmenting the prediction performance [54].
The first step of the modelling process consisted of fitting the models with the training data and following the notions described in Section 3.5.3 for each model accordingly to perform the model parameter tuning. Likewise, we performed a variable backward selection process which consists of iteratively removing such model variables that reduce the quality of the model by assessing multiple variable combinations [29]. This process consists of the model building and evaluation of multiple combinations of the considered best independent variables. This iterative process is repeated until the variable combination that better improves the model quality is reached.
Figure 3 provides a graphical representation of the seasonal decomposition process performed for the SARIMAX model with the biweekly SQALE index data of the Apache httpscore project to achieve the best model parameters. Described in a top-down structure, the first plot provides the observed progress of the SQALE index across time, the second and the third plots provide the existing trend and seasonality pattern, and the fourth one depicts the remaining residuals or noise in the data after de-trending and de-serializing has been performed. The seasonal decomposition process is conducted every time a seasonally adjusted model is built. Hence, we provide the visualization of the seasonal decomposition for each of the considered projects in the provided online appendix (see Section 3.7).
For each fitted TSA model, the explained iterative process concludes after all the combinations have been explored and the best parameter order has been found. This stage provides a model that shows a good fit for the data and provides model residuals normally distributed, thus showing good model quality. We provide a graphical representation of this stage in Figure 4, where the model diagnostics are presented for the same example project as in Figure 3.
The standardized residuals help to identify the existing anomalies or obvious patterns if the former ones are not centered around zero, therefore meaning that some patterns remain uncaptured by the model. The second plot provides the histogram of the model residuals where in comparison with the standard normal distribution the kernel density estimate seems to be approximately normally distributed. In the third plot, the model residuals are compared against the Q-Q plot standards, showing a deviated distribution of the residuals from normality. Finally, the correlogram or Auto Correlation Function (ACF) plot shows the residuals autocorrelations, which should not be after the model is fitted. Hence, they should not be exaggeratedly outside the marked confidence interval.
3.6.2 Execution process for ARIMA+LM and SARIMA + LM models
Similarly to the previous section, the ARIMA+LM and SARIMA+LM approaches share the same model-fitting logic, with the difference being the adjustment of the seasonality effect. Therefore, for the sake of not repeating model descriptions, if we consider, for instance, the SARIMA+LM approach, it fits a SARIMA univariate model for each of the independent variables, thus predicting the future values of each variable independently and only based on the past historical observations. The LM model is then fitted with the predicted values of the independent variables in order to predict future values of the Sqale index (see Section 3.5.1). Then, we performed the parameter-turning process for each of the models to achieve the best model parameter combination to be used for prediction as shown in 3.5.3. We then predicted the future values of each of the independent variables through their trained SARIMA models, respectively, following the performance evaluation methodology described in Section 3.5.6. With the future values of the independent variables already predicted, we trained the LM model with the training set of the historical SQALE index values along with the historical values of the independent variables. Thus, during the testing stage, we used the independently predicted values of the independent variables to predict the future SQALE index values. The performance evaluation of the linear regression model was computed in the same format as previously done in Section 3.6.1.
| Apache projects | Time frame⋄ | Original data | Monthly data | Biweekly data | |||||||||
| # | Min/Max SI | Mean SI | Std SI | # | Min/Max SI | Mean SI | Std SI | # | Min/Max SI | Mean SI | Std SI | ||
| archiva | 05/11-02/12 | 1,630 | 0−172,634 | 71,068 | 34,803 | 65 | 0−116,422 | 68,924 | 32,639 | 136 | 0-121,632 | 68,916 | 31,354 |
| batik | 10/00-08/02 | 941 | 65,831−175,362 | 132,111 | 29,957 | 21 | 71,725−174,077 | 138,100 | 30,581 | 44 | 71,725−174,968 | 140,551 | 29,246 |
| bcel | 10/01-04/18 | 291 | 47,854−59,608 | 54,304 | 3,262 | 173 | 47,869−58,821 | 54,348 | 3,223 | 394 | 47,869−58,821 | 54,426 | 3,233 |
| beanutils | 03/01-10/18 | 324 | 1,380−58,250 | 39,025 | 14,346 | 190 | 1,380−58,142 | 47,057 | 12,111 | 427 | 1,380−58,142 | 47,154 | 11,839 |
| cayenne | 01/07-07/08 | 353 | 179,813−223,285 | 205,411 | 12,555 | 17 | 179,813−222,828 | 203,314 | 13,481 | 33 | 179,813−222,875 | 203,402 | 13,344 |
| cocoon | 03/03-02/07 | 2,311 | 64,810−286,277 | 183,528 | 65,634 | 46 | 66,640−282,280 | 181,404 | 75,407 | 95 | 65,370−283,313 | 179,951 | 73,980 |
| codec | 05/03-05/18 | 229 | 2,078−17,707 | 10,618 | 4,567 | 164 | 2,088−17,365 | 10,216 | 4,595 | 365 | 2,078−17,707 | 10,317 | 4,553 |
| collections | 04/01-08/16 | 729 | 1,494−70,377 | 40,864 | 17,315 | 155 | 1,504−68,115 | 46,948 | 13,180 | 349 | 1,504−67,959 | 47,377 | 12,689 |
| cli* | 06/02-03/17 | 240 | 0−17,714 | 8,610 | 4,770 | 162 | 0−17,519 | 8,281 | 3,800 | 365 | 0−17,519 | 8,282 | 3,743 |
| exec* | 08/05-10/17 | 100 | 2,126−5,537 | 3,638 | 916 | 136 | 2,258−5,537 | 3,908 | 1,047 | 303 | 2,126−5,537 | 3,876 | 1,041 |
| fileupload* | 07/02-12/14 | 88 | 768−6,107 | 3,687 | 1,719 | 140 | 788−5,698 | 4,292 | 1,315 | 312 | 788−5,848 | 4,289 | 1,299 |
| io* | 01/02-11/17 | 396 | 249−42,328 | 15,782 | 8,402 | 160 | 249−42,328 | 18,970 | 10,420 | 361 | 249−42,328 | 19,134 | 10,128 |
| jelly* | 02/02-09/17 | 311 | 1,209−42,910 | 24,178 | 11,563 | 182 | 1,209−42,910 | 40,079 | 6,850 | 394 | 1,209−42,910 | 40,192 | 6,437 |
| jexl* | 04/02-01/18 | 351 | 3,338−20,048 | 12,831 | 5,562 | 164 | 4,363−19,994 | 14,537 | 4,414 | 368 | 3,567−20,040 | 14,605 | 4,423 |
| configuration | 12/03-09/18 | 764 | 4,498−44,634 | 27,754 | 10,463 | 147 | 5,182−44,634 | 29,486 | 9,959 | 323 | 4,890−44,634 | 29,923 | 9,817 |
| daemon | 09/03-11/17 | 28 | 2,290−4,168 | 3,133 | 652 | 168 | 2,290−4,062 | 3,150 | 775 | 362 | 2,290−4,062 | 3,148 | 773 |
| dbcp | 04/01-11/16 | 298 | 5,364−19,459 | 13,570 | 3,506 | 168 | 5,364−19,386 | 13,626 | 3,445 | 368 | 5,364−19,319 | 13,636 | 3,376 |
| dbutils | 11/03-04/05 | 63 | 3,357−7,954 | 4,935 | 1,339 | 129 | 3,357−6,782 | 4,671 | 1,193 | 288 | 3,357−6,782 | 4,696 | 1,192 |
| digester | 05/01-08/17 | 340 | 20−31,542 | 13,420 | 6,825 | 179 | 20−31,525 | 21,636 | 9,356 | 397 | 20−31,525 | 21,627 | 6,700 |
| felix | 08/05-07/09 | 1,149 | 22,667−283,638 | 144,966 | 73,749 | 44 | 22,667−283,530 | 121,831 | 75,758 | 88 | 24,405−265,380 | 120,187 | 74,220 |
| hive | 09/08-03/15 | 3,001 | 73,317−701,914 | 434,703 | 201,776 | 75 | 73,332−701,210 | 288,531 | 187,519 | 154 | 73,332−687,974 | 289,030 | 186,604 |
| httpclient | 12/05-05/17 | 902 | 55−85,666 | 48,009 | 25,366 | 114 | 55−85,666 | 49,809 | 26,022 | 244 | 55−85,666 | 49,205 | 25,701 |
| httpcore | 02/05-05/17 | 666 | 547−86,264 | 36,145 | 24,797 | 131 | 547−81,133 | 47,585 | 19,596 | 284 | 547−83,497 | 47,663 | 19,368 |
| jxpath | 09/01-05/18 | 148 | 23,071−32,029 | 28,448 | 2,500 | 188 | 24,708−32,029 | 29,013 | 1,840 | 422 | 23,071−32,029 | 29,032 | 1,896 |
| net | 04/02-10/18 | 543 | 12,516−34,898 | 23,536 | 3,347 | 169 | 12,706−30,433 | 24,283 | 4,163 | 374 | 12,706−30,433 | 24,372 | 4,088 |
| ognl | 05/11-06/18 | 72 | 20,410−34,077 | 27,278 | 5,615 | 85 | 20,437−34,006 | 26,774 | 1,855 | 183 | 20,444−34,006 | 26,712 | 1,717 |
| santuario | 10/01-10/18 | 914 | 2,201−126,431 | 71,472 | 25,716 | 162 | 2,201−126,431 | 80,389 | 27,220 | 371 | 2,201−126,431 | 82,370 | 27,046 |
| thrift | 05/08-10/12 | 145 | 0−21,395 | 15,965 | 3,629 | 42 | 0−21,026 | 16,486 | 4,347 | 96 | 0−21,395 | 16,757 | 4,057 |
| validator | 01/02-10/18 | 249 | 39−17,541 | 17,851 | 4,090 | 184 | 39−17,541 | 12,726 | 3,616 | 407 | 39−17,541 | 12,791 | 3,549 |
| vfs | 07/02-07/18 | 428 | 4,175−29,340 | 17,851 | 7,193 | 168 | 4,175−29,340 | 21,431 | 6,168 | 373 | 4,175−29,340 | 21,667 | 5,862 |
| zookeeper | 07/14-12/16 | 58 | 57,033−64,173 | 60,068 | 1,926 | 24 | 57,033−64,090 | 60,116 | 1,971 | 52 | 57,033−64,173 | 60,150 | 1,997 |
| #: Number of data points, SI: SQALE Index, *: Apache Commons project | |||||||||||||
| : The dates are approximations as they slightly differ between the projects in some cases. The date format is mm/yy | |||||||||||||
3.7 Replicability
To allow verifiability and replicability, we made the raw data and analysis files available in our online appendix.666 https://figshare.com/s/61eeca3d832408c9424f
4 Results
In this section, we report the results obtained by answering our RQs. The dataset included original commit data for 31 OS Java projects and the serialization and posterior linear interpolation created data serialized in biweekly and monthly time series. We provide results for 14 software projects which were successfully processed through all the adopted models in this study. We encountered linear algebra convergence issues while fitting the TSA models. We report this issue in the model fitting stage as a threat to the validity of this study, which we formally state in Section 6 and provide further details on it later in this section. Therefore, since we encountered this issue in the execution of different models across the collected projects, and for the sake of fairness in the models’ results, we excluded projects that were not able to provide results for all the considered models in the study. To answer our RQs, we present the results of the performed analysis as the average score of the outputs obtained with each prediction model for the defined prediction performance metrics. Likewise, the aggregation step is performed for the results obtained with biweekly and monthly data separately. Figure 6 provides an illustrated example of the MAE results obtained for the executed models in each of the analyzed projects. We provide all the result figures and tables before aggregation in the shared online appendix (see Section 3.7).
4.1 Descriptive Analysis
We now present the descriptive statistics of the initially considered subjects before building our prediction models. Table 2 shows the statistical characteristics of the projects within the raw data set as well as the generated time series data. From the presented table, we discovered two opposite trends in the frequency of the collected projects used to perform SQ analyses in their code. On one side, we observed projects whose number of analyzed commits in the collected raw dataset decreased during the time serialization to achieve the monthly and biweekly datasets (e.g. felix in Table 2). This phenomenon was given in projects whose SQ analysis activity was intensive during the period of time in which the data collection was performed. On the other side, some projects presented the inverse situation as their number of observations increased during the time serialization process (e.g. fileupload in Table 2). This reflected a scarce use of of SQ during the time frame of collected observations and therefore required an intensive data interpolation action to fill the missing time observations.
At this stage of the analysis, we realized that depending on the frequency of the execution of SQ analyses, the quality of the data would vary. The presented variations suggested that performing static analysis with tools like SQ in a fixed frequency could result in concise, authentic, and well-monitored time series data on the state of the quality of the code base of projects. In contrast, less frequent or randomly executed SQ analyses provide insights into the code base throughout time but leave notable periods of time with missing data, thus requiring the action of data interpolation to perform TSA. Moreover, due to the missing data for some of the projects, the algorithms from the adopted TSA models failed to converge during the model fitting stage described in Section 3.5.1. Due to the existing complexity in the data, we encountered several linear algebra decomposition issues [55] involved in the parameter estimation process while fitting the TSA models. As mentioned earlier, we leveraged preprocessing the set of independent variables within the model training through the backward variable selection criteria as well as standardization. However, Table 3 reports the cases in which the adopted models reported linear algebra decomposition issues, which impeded obtaining results for all the models in each of the projects that reported issues respectively.
| Apache projects | Time frame⋄ | Monthly data | Biweekly data | ||||||
| ARIMAX | ARIMA+LM | SARIMAX | SARIMA+LM | ARIMAX | ARIMA+LM | SARIMAX | SARIMA+LM | ||
| archiva | 05/11-02/12 | ||||||||
| batik | 10/00-08/02 | ||||||||
| bcel | 10/01-04/18 | ||||||||
| beanutils | 03/01-10/18 | ||||||||
| cayenne | 01/07-07/08 | ||||||||
| cocoon | 03/03-02/07 | ||||||||
| codec | 05/03-05/18 | ||||||||
| collections | 04/01-08/16 | ||||||||
| cli* | 06/02-03/17 | ||||||||
| exec* | 08/05-10/17 | ||||||||
| fileupload* | 07/02-12/14 | ||||||||
| io* | 01/02-11/17 | ||||||||
| jelly* | 02/02-09/17 | ||||||||
| jexl* | 04/02-01/18 | ||||||||
| configuration | 12/03-09/18 | ||||||||
| daemon | 09/03-11/17 | ||||||||
| dbcp | 04/01-11/16 | ||||||||
| dbutils | 11/03-04/05 | ||||||||
| digester | 05/01-08/17 | ||||||||
| felix | 08/05-07/09 | ||||||||
| hive | 09/08-03/15 | ||||||||
| httpclient | 12/05-05/17 | ||||||||
| httpcore | 02/05-05/17 | ||||||||
| jxpath | 09/01-05/18 | ||||||||
| net | 04/02-10/18 | ||||||||
| ognl | 05/11-06/18 | ||||||||
| santuario | 10/01-10/18 | ||||||||
| thrift | 05/08-10/12 | ||||||||
| validator | 01/02-10/18 | ||||||||
| vfs | 07/02-07/18 | ||||||||
| zookeeper | 07/14-12/16 | ||||||||
| *: Apache Commons project | |||||||||
| : The dates are approximations as they slightly differ between the projects in some cases. The date format is mm/yy | |||||||||
4.2 On the performance of Code TD prediction with existing time-dependent approaches (RQ1)
We conducted the prediction analysis described in Section 3.6 for each of the 31 projects in our dataset. Figure 5 shows the unified graphical illustration of the predictive performance results from the adopted performance metrics. The displayed results demonstrate a clear superiority of ARIMAX compared to ARIMA+LM with both biweekly and monthly serialized time series data. Considering the results for MAE and RMSE, the differences in prediction error between the ARIMAX and ARIMA+LM are 14,000 and 14,666 within the biweekly data, and 15,237 and 15,931 within the monthly data, thus denoting a remarkable difference among both approaches. Hence, we can affirm that given the existing TSA models for Code TD prediction, ARIMAX model overcomes the predictive performance of ARIMA+LM model. To be concise, both models performed better when being trained with biweekly data as depicted by MAE and RMSE, whereas MAPE results provided an improvement in the predictive performance for ARIMA+LM model when training it with monthly data.
4.3 On the comparison of time-dependent approaches and ML models for Code TD prediction (RQ2)
We compared the ARIMAX and ARIMA+LM models against the adopted ML algorithms commonly used in the field. Tables 4 and 5 present the prediction performance results for the biweekly and monthly data, respectively. More specifically, the tables provide the average results calculated from the prediction results obtained for the 14 projects. Results for the individual projects are included in the replication package. It is worth noting that in both cases, using biweekly and monthly data, ARIMAX model demonstrated a clear superiority over the rest of the models, thus providing a more accurate prediction outcome for future Code TD observations. Results such as 1,525.96 for MAE, as well as 2,646.20 for RMSE within the biweekly results, compared to the ones presented by the next model following the ranking of results, RF, denote a quantified improvement of 717 and 1,232 in the described prediction errors accordingly. Moreover, RF proved to be the second-best prediction model. As described previously, RF assumes non-linear relationships between variables. Consequently, despite the indications that TSA models such as ARIMAX may be appropriate for predicting Code TD, the potential for non-linear models to capture intricate relationships cannot be discounted. This is particularly relevant in light of the possibility of incorporating further data in the future and additional model predictors, for instance. Looking at the rest of the ML models and excluding the special case of SVM, non-linear models present more accurate results than linear models in all the calculated metrics.
| Approach | MAPE (%) | MAE | RMSE |
| SARIMAX | 1.00 | 1,539.92 | 2,721.95 |
| SARIMA + LM | 17.67 | 12,921.79 | 14,412.90 |
| ARIMAX | 0.97 | 1,525.96 | 2,646.20 |
| ARIMA + LM | 17.35 | 15,525.21 | 17,312.14 |
| MLR(L) | 4.85 | 4,895.50 | 6,759.43 |
| SVM(NL) | 23.87 | 38,929.00 | 39,904.36 |
| XGB(NL) | 1.72 | 3,006.66 | 4,686.55 |
| RF(NL) | 1.44 | 2,242.57 | 3,878.84 |
| SGD(L) | 5.04 | 5,494.12 | 7,275.05 |
| L1(L) | 5.18 | 5,431.16 | 7,112.04 |
| L2(L) | 4.86 | 4,898.60 | 6,709.81 |
| Approach | MAPE (%) | MAE | RMSE |
| SARIMAX | 1.41 | 2,078.64 | 3,145.27 |
| SARIMA + LM | 13.32 | 13,871.48 | 15,842.22 |
| ARIMAX | 1.65 | 2,349.81 | 3,469.03 |
| ARIMA + LM | 15.02 | 17,586.97 | 19,400.41 |
| MLR(L) | 4.62 | 3,978.65 | 5,567.30 |
| SVM(NL) | 23.12 | 27,822.99 | 28,620.02 |
| XGB(NL) | 5.80 | 4,600.92 | 5,950.18 |
| RF(NL) | 1.90 | 2,572.85 | 4,202.80 |
| SGD(L) | 4.89 | 4,478.40 | 5,956.63 |
| L1(L) | 5.09 | 4,609.15 | 6,067.31 |
| L2(L) | 4.50 | 3,706.95 | 5,192.83 |
4.4 On the impact of seasonality on Code TD prediction in time-dependent approaches (RQ3)
Following the set of defined research questions, we wanted to analyze the impact of addressing the seasonality effect on the prediction performance of TSA models through our RQ3. Thus, we conducted the prediction analysis described in Section 3.6 with SARIMAX and SARIMA+LM models. The process was similar to RQ1 and used the same data for this RQ. Figure 5 shows the unified graphical depiction of the predictive performance results from the adopted performance metrics. To avoid overwhelming the reader with too many plots, we displayed the results for the seasonally adjusted models along with their former counterpart models ARIMAX and ARIMA+LM. The displayed results demonstrate a systematic difference in prediction performance among the multivariate single model approach, i.e. ARIMAX and SARIMAX, compared to that of the multivariate combined model approach, i.e. ARIMA+LM and SARIMA+LM. Therefore, we can affirm that the superiority in prediction performance demonstrated on ARIMAX compared to ARIMA+LM is maintained with SARIMAX and SARIMA+LM when adjusting the models for capturing seasonality. Figure 5 demonstrates a clear superiority of SARIMAX over SARIMA+LM given biweekly as well as monthly serialized time series data. Hence, we can affirm that given the adjustment of the existing TSA models for Code TD prediction, the SARIMAX model overcomes the predictive performance of the SARIMA+LM model. Precisely, considering for instance the results for MAE and RMSE the difference regarding the prediction error between the SARIMAX and SARIMA+LM stand in 11,382 and 11,691 within the biweekly data, and 11,792 and 12,697 within the monthly data respectively, thus denoting a remarkable difference among both approaches. Answering to RQ3, the displayed results provide slightly similar results between SARIMAX and ARIMAX models. Results for biweekly data provided better prediction performance for ARIMAX compared to SARIMAX while the opposite happened for monthly data. MAPE results provided worse prediction performance for SARIMA+LM than for ARIMA+LM when using biweekly data. However, SARIMA+LM overtook ARIMA+LM for the MAPE results when using monthly data, as well as in the other prediction performance evaluation metric results. Therefore, we can affirm that adjusting the TSA models to capture the seasonality effects of the data improves their prediction performance to some extent.
4.5 On the comparison of seasonally adjusted time-dependent models and ML models for Code TD prediction (RQ4)
Finally, we compared the previously confronted SARIMAX and SARIMA+LM models against the adopted ML algorithms commonly used in the field. Tables 4 and 5 present the prediction performance results for SARIMAX and SARIMA+LM models compared to the adopted ML models. As expected from the results obtained for RQ3, where SARIMAX provided similar prediction results as ARIMAX when using biweekly data and overtook ARIMAX when using monthly data, in this case, the SARIMAX model also demonstrated a clear superiority among the rest of the compared models, thus providing a more accurate prediction outcome for future Code TD observations. It is worth noticing that capturing the seasonality effect did not change the key results observed for RQ2. In fact, SARIMAX did not improve the results of ARIMAX model using biweekly data, and the improvement experimented compared to ARIMAX with monthly resulted to be low. Moreover, RF demonstrated to be the second-best prediction model. Therefore, even if the results suggest the suitability for linear models, non-linear models may also offer the potential for capturing more complex relationships. Looking at the rest of the ML models and excluding the special case of the SVM model, non-linear models present more accurate results than linear models in all the calculated metrics.
5 Discussion
The objective of the presented study was to analyse the impact of temporal factors in Code TD prediction. Code TD was approximated by leveraging the SQALE index metric calculated by SQ, and the predictions were made including the set of different types of code smell issues in the projects within the prediction models. The adoption of TSA prediction models enabled the empirical demonstration of the suitability of time-dependent techniques for Code TD prediction, as well as a comparative analysis with commonly used ML prediction algorithms. Consequently, 11 distinct prediction models were constructed for each project. The analysis included a comprehensive data collection, preprocessing, and examination of 31 open-source Java projects. However, in order to ensure the integrity and fairness of the results, only the outcomes for the 14 projects that provided complete results with the adopted prediction models were considered.
The study was divided into four research questions. The first research question compared the performance of two ARIMA-founded time-dependent approaches already implemented in the academic literature. Following our initial assumption, the ARIMAX model previously proposed by by Mathioudaki et al. [11] outperformed the combination model ARIMA+LM proposed by Zozas et al. [12]. Thus, our initial takeaway depicted the superiority of ARIMAX over ARIMA+LM model when predicting Code TD. These results suggest that, given a time-dependent prediction context, multivariate models which include and simultaneously predict the independent variables as well within the prediction model such as ARIMAX succeed on providing robust results as shown in Figure 5.
Furthermore, in the second research question we proved the robustness reported with ARIMAX by comparing its Code TD prediction performance against a set of the most common ML algorithms adopted in the SE field. The results provided in Tables 5 and 4 corroborated this evidence. These results suggest the necessity of adopting models that consider the time dependence factor in the dependent variable to be predicted. However, it is worth noticing the close results reported by models such as Random Forest or Extreme Gradient Boost, prediction algorithms widely used in academia as well as in industry. Given the non-linear distributional assumption of these models, and the oppositely linear distributional assumption, which is the foundation for TSA models such as ARIMAX, leveraging the combination of these models through ensemble learning approaches might further improve the prediction performance reported in this study.
Moreover, we analyzed the impact of addressing the seasonal component within the time series data on the obtained results. For that, we leveraged the model extensions provided in ARIMA-founded models to analyze the results for SARIMAX and SARIM+LM models. As the takeaway of the performed analysis, we observed slight improvements in the initially adopted TSA models, thus depicting a small impact on the prediction performance from controlling the seasonality component. This evidence suggests a small dependence on seasonality from the perspective of the OS software development community, which might not follow structured development schedules as compared to the software development activity performed in the industry.
As the key takeaway of this study, we can confirm that time-dependent models are applicable and competitive in TD prediction, not only at a single project level but also at an aggregated level as we have demonstrated. Further research is still needed to understand the methodology required to prepare the most suitable data for TSA models. It has been demonstrated that if the data fits well in the TSA model, this latter one can provide high prediction performance. Therefore, this study suggests the adoption of periodic software quality analysis practices within the project development process, thus providing a thorough monitorization of metrics such Code TD and better quality data for prediction performance improvement. Similarly, it is important to note at this stage that when we make our predictions, these values are the obtained most likely estimates, therefore we are not comparing our real values against the result of a simulation or an execution.
TSA models can have a considerable prediction performance, but there is a need for further research in modeling TD prediction while considering time as a relevant factor.
6 Threats to Validity
Construct validity
Some projects have been affected by only a few code smell rule issues, so we decided to exclude them as their independent variables were considered uninformative. Additionally, the irregularity of the commit data can cause missing data for the generated periodical time series. The threat was minimized by linearly interpolating values for the missing periods. Approximations cannot be as accurate as using real data, therefore we could expect different results if periodically complete real data existed.
Internal validity
We selected one of the metrics that depict technical debt, i.e. SQALE index as the variable of interest, and the set of code smell types that SQ reports as the reason for the given SQALE index. However, we did not consider collecting additional independent variables that might affect the results. Therefore, we could expect different results if another choice of independent variables had been made.
External validity
The study subjects were mature open source projects written in Java and fulfilling the rigorous quality standards of ASF. The included projects represent a wide spectrum of application domains, including external libraries, frameworks, web utilities, and substantial computational infrastructures. Hence, the obtained results can be generalized to projects with similar characteristics, and therefore non-mature projects, retired projects, and projects not using SonarQube are excluded.
Conclusion validity
This study applied commonly known statistical and ML techniques. During the data analysis, we ensured the assumptions of the techniques were fulfilled. However, a low number of data points in some projects may reduce the prediction power of the used models. Additionally, the ARIMA + LM model may not capture the time-dependent nature of the SQALE index, as its prediction is performed through a Multiple Linear Regression model, which does not consider the time dependence factor. Regarding the reliability of the measures, the measurements of the variables were collected from the dataset, and the construction of the final set of model variables was made through an automated process. For the sake of reliability on the results, we excluded the collected projects which did not allow the algorithms from the adopted models to converge within the model fitting. Therefore, we report only the aggregated results from the projects which provided results for the entire set of considered models in this study.
7 Related Work
In this section, we report the related work including the current status of Technical Debt (TD) Prediction.
Tsoukalas et al. [56] investigated the effectiveness of ML techniques in modeling and predicting the evolution of TD. Specifically, they carried out an extensive investigation using a dataset they compiled, consisting of weekly snapshots of fifteen open-source software projects spanning three years. This dataset is augmented with software-related metrics extracted from two commonly used static analysis tools, which serve as predictors for TD. Using these identified predictors, they develop a range of TD predicting models employing well-known ML algorithms. These models were then tested against various timeframes for prediction. The findings from their analysis revealed that linear regularization models aptly capture and yield meaningful predictions for TD evolution over short prediction periods. On the other hand, non-linear random forest regression outperforms linear models over longer prediction horizons. In a second work, Tsoukalas et al. [9] adopted ML models to classify TD classes (High/Not-High TD), analyzing 25 open source projects. The achieved results had an accuracy estimation of approximately 0.79, with a recall of 0.86.
Mathioudaki et al. [44] explored the application of Deep Learning to enhance the precision of long-term TD prediction. To achieve this, they created, assessed, and juxtaposed Deep Learning models employing a dataset encompassing five prominent real-world software applications sourced from the Technical Debt dataset version 1 [24]. Results revealed that Deep Learning techniques yield TD prediction models that display commendable predictive accuracy, extending up to 150 steps ahead in future prediction.
Aversano et al. [45] investigated the potential utility of software system quality metrics in accurately predicting TD. To achieve this, information concerning quality metrics from 8 distinct open-source software systems was examined. These metrics were then fed into various ML algorithms to predict technical debt. The outcomes demonstrated strong predictive performance, and the suggested approach offers a valuable method for comprehending the practical aspects of the technical debt phenomenon.
Focusing on the works that adopted TSA models to predict TD, we concentrated on two papers. Mathioudaki et al. [11] investigated the predictive capabilities of TSA models for TD prediction. They aimed to determine whether incorporating independent variables as predictors, known as TD predictors, into ARIMAX models [23, p. 451] could yield more precise TD predictions compared to conventional Autoregressive Integrated Moving Average (ARIMA) models. Their investigation was conducted using five datasets that captured the historical evolution of software metrics derived from static code analysis across five long-standing projects. They generated predictions using both ARIMA and ARIMAX models for various time horizons on this data. Results yielded a clear conclusion. Across the examined open-source software projects, the accuracy of ARIMAX models exceeded that of ARIMA models by a significant margin.
More recently, Zozas et al. [12] combined two models for predicting Technical Debt Principal: Supervised Backwards Stepwise Regression, and ARIMA. Based on the application of Backwards Stepwise Regression they obtained the most significant predictors to describe the TD response variable. They performed univariate TSA prediction through ARIMA for each of the selected predictors to predict future values, and then they used the previously developed regression model to estimate the future values of the response variable.
Comparing our work with Mathioudaki et al. [11], they considered the same data set we did, but they only included 5 projects out of 33, while we investigated the entire dataset. Moreover, in our work, instead of conducting a comparison between univariate and multivariate TSA, we aim to directly perform a comparison between different multivariate TSA approaches to investigate which approach obtains better prediction performance. The selected predictors will be used to predict the value of the SQALE index, which will be our target value. Our approach relies on performing different multivariate applications of the ARIMA model, and for testing its performance we aim to split the data for training and testing the model through the Walk-Forward Train-Test approach.
Furthermore, as we believe that ML models should not be trained with just one feature but the exact amount for the sake of the quality of that model, we compare in this case the performance of the mentioned multivariate TSA models against a set of chosen ML models based on previous research.
8 Conclusions
We conducted an empirical study to compare two TSA approaches, their respecting version adjusted to treat the seasonality component and seven ML algorithms for Code TD prediction as SQALE index computed by SQ.
The compared methods were trained using Code TD observation data that was serialized into biweekly and monthly periodicity levels from 14 open-source Java projects. The comparison denoted a clear superiority of the ARIMAX model compared to the rest of the adopted models when training the models with biweekly time series, while SARIMAX provided better results with monthly time series. However, results provided by SARIMAX suggested that the impact from seasonality might not improve substantially the predictive performance of ARIMAX, thus demonstrating that the main improvement between ML algorithms’ performance and TSA models is generated by the ARIMAX model. Therefore, our findings indicate that, while the seasonality factor might not strongly impact the predictive performance of time-dependent prediction techniques, considering the time-dependence factor improves the prediction results compared to that of time-agnostic techniques. This and the required data serialization performed in this study in order to analyse time series data suggest the potential benefit performing sequential software quality analyses in software projects can have in order to improve the Code TD prediction performance.
In future work, we aim to analyze the implementation of transformer-driven prediction techniques. The adaptation of TSA techniques with these models provides great modelling ability for long-range dependencies and interactions in sequential data and TSA models. This will further contribute to the research community with advanced techniques for predicting and therefore anticipating Code TD.
Appendix: Parameter tuning of the Box-Jenkings models
In this appendix, we describe the mathematical definitions of the model parameters composing the TSA models included in the work. The adopted models stand as modifications of the Box-Jenking model, better known as ARIMA [34]. This model is a univariate TSA model composed of the autoregressive component, moving average component, and differencing component, as explained below.
8.1 Autoregression (AR) order (p / P parameter)
The autoregressive process of order p, AR(p), denotes a stationary time series such that
where are fixed autoregression coefficients and are independent random noises of mean and constant variance . An AR(p) process is stationary if the p roots of fall within the random state unit circle [34].
Seasonality adjustment occurs when the autoregressive pattern occurs at multiple lags of the defined periodicity (m). Thus a seasonal autoregression (SAR) process is defined by
where is defined as .
8.2 Moving Average (MA) order (q / Q parameter)
The moving average process of order q, MA(q) denotes a stationary time series such that
where and is defined as . Every MA(q) process is stationary, but there can be two different MA(q) processes with the same autocovariance [57]. We avoid this issue by enforcing invertible MA processes (see details in Section 3.7).
As before, a seasonality adjustment is performed when the moving average pattern occurs at multiple lags of the defined periodicity (m). Thus seasonal moving average (SMA) processes are defined by
where is defined as .
8.3 Differecing order (d / D parameter)
Among the existing approaches to reach stationarity, models such as ARIMA apply differencing repeatedly to the series of data until the differenced observations resemble a realization of stationary time series, required for making accurate forecasting [57]. Considering for instance an original time series we can define the first differencing order (d = 1) as:
Hence, the differencing operator or lag-d can be formulated as
Moreover, differencing can also be used in cases where the seasonality of the series is addressed. If the series has a seasonal pattern, then d-lag differencing order can remove this pattern. Thus, if and had period d, then:
where stands as the seasonality component at series observation t, is the trend components accordingly, and is the existing random noise.